OpenStudy (anonymous):

Explain how raising a quotient to a power is similar to raising a product to a power

OpenStudy (amistre64):

a quotient is a product ...

OpenStudy (amistre64):

\[\frac ab=\frac a1\frac 1b\]

OpenStudy (dan815):